# Dynamic Gaussians Mesh: Consistent Mesh Reconstruction from Monocular Videos

### v1

@@ Abstract
Project page: https://www.liuisabella.com/DG-Mesh/ Equal advisory
@@ 1 Introduction
Our key observation is that the direct training of 3D Gaussians mentioned above will lead to uneven spreading of Gaussians in 3D space (as shown in Fig. 2 left). However, this is detrimental for mesh reconstruction performance due to the violation of Possion solver assumption that points should be evenly distributed; worse still, non-uniform Gaussians make it trickier to track across a long time period. Thus, we propose the Gaussian-Mesh Anchoring procedure during training, which yields uniformly distributed Gaussians at each frame. Concretely, first, we encourage deformed 3D Gaussians in each time step to align with the corresponding mesh surface that has uniform face distribution due to the underlying iso-surface algorithm; and second, we allow adding Gaussians in the world frame at each time step if certain mesh faces are not covered by existing deformed Gaussians, as well as deduplicting Gaussians for mesh faces that are covered multiple times. We call the new set of Gaussians with 1-1 correspondences to mesh faces as anchored Gaussians. As depicted in Fig. 2, anchored Gaussians spread more uniformly in 3D.
@@ 3 Method
Our method maintains a set of canonical 3D Gaussians and deformation networks to deform the Gaussians into different times (Sec. 3.1). We use a differentiable Poison Solver and a differentiable Marching Cubes method to reconstruct the oriented 3D Gaussian points into meshes (Sec. 3.2). Our model allows the mesh-to-mesh correspondence. This is constructed by two types of correspondence, including mesh-to-point correspondence under each time step and the points to canonical points correspondence across all time frames. We propose the Gaussian-Mesh Anchoring to encourage the 1-1 correspondence between the mesh faces and the Gaussians, producing more uniformly distributed anchored Gaussians. The Cycle-Consistent Deformation maintains the points to canonical points correspondence, which includes a backward deformation network that computes the deformation from the anchored Gaussians to the canonical Gaussians (Sec. 3.3). Finally, we combine all the training objectives (Sec. 3.4).
@@ 3.2 Mesh Reconstruction
Laplacian Regularization. In order to better preserve the mesh surface tessellation, we adopt a mesh Laplacian regularization term to penalize local curvature changes. We use a uniformly weighted differential $\delta_{i}$ of vertex $v_{i}$ to describe the difference between the position of vertex $v_{i}$ and the average position of its neighbors. The Laplacian regularization term is the mean square $\delta_{i}$ for all vertices in the mesh, and $n$ is the total number of vertices of the mesh.
@@ 3.3 Dynamic Mesh Correspondence
Gaussian-Mesh Anchoring. To build the mesh-to-point correspondence at each time frame, we introduce the Gaussian-Mesh Anchoring. The original 3D Gaussian densification and deformation mechanism leads to uneven spreading of Gaussians in 3D space (Fig. 2), which is undesirable for finding the 1-1 correspondence between mesh and the Gaussians. On the other hand, DPSR discretizes the function values and differential operators in the Poisson equation by assuming the normal vector field $n$ and indicator function $\chi$ are uniformly sampled along each dimension. However, this uniformity is not guaranteed during the above training procedure. To encourage the uniformity of the 3D Gaussian distribution, we use reconstructed mesh to guide the densification and pruning of the 3D Gaussian in the deformed space. Specifically, given a deformed 3D Gaussian set $\{G(x_{i})\}_{i=0}^{N}$ and the set of the face cetroids from its reconstructed mesh $\{f_{j}\}_{j=0}^{M}$ and $N\neq M$. For each 3D Gaussian point $x_{i}$ we find its nearest neighbor in the mesh face set:
@@
During the densification and pruning process, for face $f_{j}$ that is the nearest neighbor to multiple $K$ Gaussian points, we prune the original $K$ Gaussians and create a new Gaussian by averaging the properties from the $K$ Gaussians:
@@
where $n$ is the total number of faces that have one-to-one correspondence to the 3D Gaussian points. The obtained anchored Gaussians $G^{\prime}$ are applied in the Cycle-Consistent deformation loss $\mathcal{L}_{cycle}$ which will be introduced in the next paragraph.
Cycle-Consistent Deformation. Deformable 3D Gaussians provide explicit motion and correspondence across times because of their point-cloud representation. To transfer the correspondence from moving Gaussian points to the dynamic mesh sequence, we align the reconstructed mesh faces and the 3D Gaussian points in the deformed space during Gaussian-Mesh Anchoring. To build the points to canonical point correspondence and update the canonical space to accommodate the anchored Gaussians in the deformed space, we encourage the cycle consistency between the canonical and deformed spaces to ensure that the anchoring performed under the deformed space can be applied back to the canonical space. We define a forward transformation function $\mathcal{F}_{f}$ and a backward transformation function $\mathcal{F}_{b}$. Given a canonical 3D Gaussian $G$ and its new form $G^{\prime}$ under deformed space, we have:
@@ 3.4 Combining All Training Objectives
We use rendering loss from both mesh rasterization and 3DGS to optimize the geometry and appearance. Given the ground truth image $I_{gt}$ and the Gaussian splatted image $I_{gs}$. The image loss term of 3DGS can be described as:
@@
Same for the rasterized mesh image loss $\mathcal{L}_{mesh}$. We also compute the $L_{1}$ loss on rasterized mask $\mathcal{L}_{mask}$.
To build the mesh-to-mesh correspondence at different time frames, the $\mathcal{L}_{anchor}$ described in Sec. 3.3 encourage the 1-1 mapping between the mesh faces and Gaussians and produce the mesh-to-point correspondence. The $\mathcal{L}_{cycle}$ encourages the points to canonical points correspondence. The final loss term can be described as:
$\mathcal{L}=\mathcal{L}_{gs}+\mathcal{L}_{mesh}+\mathcal{L}_{mask}+\mathcal{L}_{lap}+\mathcal{L}_{anchor}+\mathcal{L}_{cycle}$
@@ 4.4 Ablation Study
Figure 7: We provide the result comparison of our method with other baselines with our dataset that has mesh ground truth. Our method gives better mesh rendering quality as well as geometry. For hard targets such as the flying bird, other methods fail to recover the thin geometry of the bird wings. Instead, our method produces satisfying results in these areas.

### v2

@@ Abstract
footnotetext: Project page: https://www.liuisabella.com/DG-Mesh/footnotetext: Equal advisory
@@ 1 Introduction
We achieve this by decomposing the mesh-to-mesh correspondence into two separate correspondences that are easier to acquire: the mesh-to-points correspondence that maps the mesh faces to the Gaussian points in each time frame; and the points-to-canonical-points correspondence that helps tracking the movement of mesh across time. Our key observation is that the direct training of 3D Gaussians mentioned above will lead to uneven spreading of Gaussians in 3D space (as shown in Fig. 2 left). However, this is detrimental for mesh reconstruction performance due to the violation of the Possion solver assumption that points should be evenly distributed; worse still, non-uniform Gaussians make it trickier to track across a long time period. Thus, we propose the Gaussian-Mesh Anchoring procedure during training, which yields uniformly distributed Gaussians at each frame. Concretely, first, we encourage deformed 3D Gaussians in each time step to align with the corresponding mesh surface that has uniform face distribution due to the underlying iso-surface algorithm; and second, we allow adding Gaussians in the world frame at each time step if certain mesh faces are not covered by existing deformed Gaussians, as well as deduplicting Gaussians for mesh faces that are covered multiple times. We call the new set of Gaussians with 1-1 correspondences to mesh faces as anchored Gaussians. As depicted in Fig. 2, anchored Gaussians spread more uniformly in 3D.
@@ 3 Method
Our method maintains a set of canonical 3D Gaussians and deformation networks to deform the Gaussians into different times (Sec. 3.1). We use a differentiable Poison Solver and a differentiable Marching Cubes method to reconstruct the oriented 3D Gaussian points into meshes (Sec. 3.2). Our model allows the mesh-to-mesh correspondence. This is constructed by two types of correspondence, including the mesh-to-point correspondence under each time step and the point-to-canonical-point correspondence across all time frames. We propose the Gaussian-Mesh Anchoring to encourage the 1-1 correspondence between the mesh faces and the Gaussians, producing more uniformly distributed anchored Gaussians. The Cycle-Consistent Deformation maintains the points to canonical points correspondence, which includes a backward deformation network that computes the deformation from the anchored Gaussians to the canonical Gaussians (Sec. 3.3). Finally, we combine all the training objectives (Sec. 3.4).
@@ 3.2 Mesh Reconstruction
Laplacian Regularization. To better preserve the mesh surface tessellation, we adopt a mesh Laplacian regularization term to penalize local curvature changes. We use a uniformly weighted differential $\delta_{i}$ of vertex $v_{i}$ to describe the difference between the position of vertex $v_{i}$ and the average position of its neighbors. The Laplacian regularization term is the mean square $\delta_{i}$ for all vertices in the mesh, and $n$ is the total number of vertices of the mesh.
@@ 3.3 Dynamic Mesh Correspondence
We construct our mesh-to-mesh correspondence with two separate correspondences: the mesh-to-points correspondence which maps the mesh faces to the Gaussian points in each time frame, and the point-to-canonical-point correspondence across all time frames that helps tracking the movement of mesh over time.
Gaussian-Mesh Anchoring. To build the mesh-to-point correspondence at each time frame, we introduce the Gaussian-Mesh Anchoring. The original 3D Gaussian densification and deformation mechanism leads to uneven spreading of Gaussians in 3D space (Fig. 2), which is undesirable for finding the 1-1 correspondence between mesh and the Gaussians. On the other hand, DPSR discretizes the function values and differential operators in the Poisson equation by assuming the normal vector field $n$ and indicator function $\chi$ are uniformly sampled along each dimension. However, this uniformity is not guaranteed during the above training procedure. To encourage the uniformity of the 3D Gaussian distribution, we use reconstructed mesh to guide the densification and pruning of the 3D Gaussian in the deformed space. Specifically, given a deformed 3D Gaussian set $\{G(x_{i})\}_{i=0}^{N}$ and the set of the face centroids from its reconstructed mesh $\{f_{j}\}_{j=0}^{M}$ and $N\neq M$. For each 3D Gaussian point $x_{i}$ we find its nearest neighbor in the mesh face set:
@@
During the densification and pruning process, as shown in Fig. 4, for face $f_{j}$ that is the nearest neighbor to multiple $K$ Gaussian points, we prune the original $K$ Gaussians and create a new Gaussian by averaging the properties from the $K$ Gaussians:
@@
where $n$ is the total number of faces that have a one-to-one correspondence to the 3D Gaussian points. The obtained anchored Gaussians $G^{\prime}$ are applied in the Cycle-Consistent deformation loss $\mathcal{L}_{cycle}$ which will be introduced in the next paragraph.
Cycle-Consistent Deformation. Deformable 3D Gaussians provide explicit motion and correspondence across times because of their point-cloud representation. During Gaussian-Mesh Anchoring, we align the reconstructed mesh faces and the 3D Gaussian points in each time frame and perform mesh-guided densification and pruning. To update the canonical space to accommodate the anchored Gaussians and build the point-to-canonical-point correspondence, we encourage the cycle consistency between the canonical and deformed spaces to ensure that the anchoring performed under the deformed space can be applied back to the canonical space. We define a forward transformation function $\mathcal{F}_{f}$ and a backward transformation function $\mathcal{F}_{b}$. Given a canonical 3D Gaussian $G$ and its new form $G^{\prime}$ under deformed space, we have:
@@ 3.4 Combining All Training Objectives
We use rendering loss from both mesh rasterization (“Mesh Image” in Fig. 3) and the 3DGS (“Gaussian Image” in Fig. 3) to optimize the mesh geometry and appearance. Given the ground truth image $I_{gt}$ and the Gaussian splatted image $I_{gs}$. The image loss term of 3DGS can be described as:
@@
Same for the rasterized mesh image loss $\mathcal{L}_{mesh}$. We also compute the $L_{1}$ loss on the rasterized mask $\mathcal{L}_{mask}$.
The Laplacian loss term $\mathcal{L}_{lap}$ in Sec. 3.2 helps preserve the mesh surface tessellation and produces smoother surface. The anchoring loss term $\mathcal{L}_{anchor}$ and the cycle-consistency loss term $\mathcal{L}_{cycle}$ described in Sec. 3.3 help the optimization of cross-frame mesh-to-mesh correspondence. The final loss term can be described as:
$\mathcal{L}=$ $\mathcal{L}_{gs}+\mathcal{L}_{mesh}+\mathcal{L}_{mask}$
$+\mathcal{L}_{lap}+\mathcal{L}_{anchor}+\mathcal{L}_{cycle}$
@@ 4.4 Ablation Study
Figure 7: We provide the reconstructed mesh visualization and mesh rendering image of our method on real-world datasets including the Nerfies dataset with monocular videos of daily life objects captured using hand-held cameras; the Unbiased4D dataset contains one sequence of real monocular video captures of a moving toy.